An oil- fired boiler takes in feed water at 75°C and produces wet steam at a pressure of 5 bars. The steam flow rate 1.50 tons/hr with a dryness fraction of 0.89. The fuel consumption rate is 6.10kg/min and its net calorific value is 41MJ/Kg. Determine the thermal efficiency of the boiler.
1
Expert's answer
2018-08-16T10:15:08-0400
Finding enthalpy of feed water, h1 and output steam h2
h1 = hf at 75 oC
h1 = 313.93 kJ/Kg
h2 = hsup at 5 bar
h2 = hf + xhfg
h2 = 640 kJ/Kg + 0.89 x 2108 kJ/Kg = 2516.12 kJ/Kg
Finding steam fl ow rate in kg/s
ms = (1.5 x 103) / 602 = 0.42 kg/s
Finding power rating of boiler.
Boiler power rating = (h2 - h1 ) x ms
Boiler power rating = (2516.12 kJ/Kg 2 - 313.93 kJ/Kg ) x 0.42 kg/s = 924.9 kJ = 0.9 MJ
Finding fuel consumption in kg/s
mf = 0.102 kg/s
Finding thermal effi ciency of boiler
η = ((h2 - h1 ) x ms) / mfCV
η = 0.9 MJ / (0.102 kg/s x 41MJ/Kg) = 0.215 = 21. 5 %